OpenStudy (anonymous):

Integrate the following [4*csc^2(3x)*cot(3x)]dx

OpenStudy (anonymous):

\[4\int\csc^23x\cot 3x~dx\] Let \(u=\cot3x\), so \(-\dfrac{1}{3}du=\csc^23x~dx\): \[-\frac{4}{3}\int u~du\]